Some​ agreed-upon procedures related to generating estimates for key variables in equation​ (2-1) follow.

a. The current​ 3-month Treasury bill rate is 3.05 ​percent, the​ 30-year Treasury bond rate is 5.32 ​percent, the​ 30-year Aaa-rated corporate bond rate is 6.74 ​percent, and the inflation rate is 2.33 percent

. b. The real​ risk-free rate of interest is the difference between the calculated average yield on​ 3-month Treasury bills and the inflation rate.

c. The​ default-risk premium is estimated by the difference between the average yields on​ Aaa-rated bonds and​ 30-year Treasury bonds.

d. The​ maturity-risk premium is estimated by the difference between the average yields on​ 30-year Treasury bonds and​ 3-month Treasury bills.

e. SanBlas​ Jewels' bonds will be traded on the New York Bond​ Exchange, so the​ liquidity-risk premium will be slight. It will be greater than​ zero, however, because the secondary market for the​ firm's bonds is more uncertain than that of some other jewelry sellers. It is estimated at 4 basis points. A basis point is one​ one-hundredth of 1 percent.

Now place your output into the format of equation​ (2-1) so that the nominal interest rate can be estimated and the size of each variable can also be inspected for reasonableness and discussion with the CFO.

What is the real​ risk-free interest​ rate?

What is the​ default-risk premium?

What is the​ maturity-risk premium?
What is the​ liquidity-risk premium?
What is the nominal interest​ rate?
Answer #1

Real Risk-Free Rate = Average Yield on 3-month Treasury Bill - Inflation = 3.05 - 2.33 = 0.72 %

Default Risk Premium = Average Yield of Aaa - rated bonds - 30 Treasury Bonds = 6.74 - 5.32 = 1.42 %

Maturity Risk Premium = Average Yield o 30 year Treasury Bond - 3 Month Treasury Bill = 5.32 - 3.05 = 2.27 %

Liquidity Premium = 4 basis points or 0.04 %

Nominal Interest Rate = Real Risk Free Rate + Inflation Rate + Default RIsk Premium + Maturity RIsk Premium + Liduidity Premium = 0.72 + 2.33 + 1.42 + 2.27 + 0.04 = 6.78 %